Find the probability that a poker hand of 5 cards will contain only black cards if it is known to contain 3 black cards.
Answer by Edwin McCravy(20054)

Since you are given that the poker hand contains 3 black cards, to be successful,
the other 2 cards are black and selected from a 49-card deck that contains 23
black cards and 26 red cards.

numerator = 23 black cards choose 2 = C(23,2) = [(23)(22)]/[(2)(1)] = 253
denominator = 49 cards choose 2 = C(49,2) = [(49)(48)]/[(2)(1)] = 1176

Answer = 253/1176

Edwin